OpenStudy (anonymous):

sound wave x has intensity 10^12 times greater than that of sound wave Y. By how much is the amplitude of x greater than the amplitude of Y?

OpenStudy (anonymous):

Now I know Intensity is proportional to the square of amplitude but I don't know how to apply that

OpenStudy (sidsiddhartha):

use \[I \alpha A^2\rightarrow \frac{ I_1 }{ I_2 }=\frac{ A_1^2 }{ A_2^2 }\]

OpenStudy (sidsiddhartha):

where I=intensity and A=amplitude

OpenStudy (anonymous):

I have never seen that notation...

OpenStudy (anonymous):

\[A _{1}^{2}\]

OpenStudy (anonymous):

Is not something I've been introduced to

OpenStudy (sidsiddhartha):

Intensity is proportional to the square of amplitude u have to use this relation

OpenStudy (anonymous):

Oh right of course so I = A^2 so A = I^0.5 giving 10^(12*0.5) giving A = 10^6 Thanks!